# Quantum channel tomography and estimation by local test

### v1

@@ 1.1 Our results
Theorem 1.1 can be viewed as an extension of the previous result by [CWZ24], which studies the power of local test for quantum states and shows that access to purifications does not help for mixed-state testing. Related results trace back to [SW22, Theorem 35], and were recently strengthened in an algorithmic sense by [TWZ25], which explicitly constructs an algorithm for generating random purifications of a mixed state. This is further leveraged in [PSTW25] for mixed-state tomography. Finally, we note that Theorem 1.1 partially answers a conjecture from [TWZ25] asserting that access to channel dilations does not help.
@@
Notably, when $rd_{2}=d_{1}$ (i.e., when the quantum channel $\mathcal{E}$ can be obtained from a unitary channel followed by tracing out a subsystem), we can achieve the Heisenberg scaling $O(1/\varepsilon)$, even if $\mathcal{E}$ is not a unitary channel.
@@
Corollary 1.3 is obtained by applying Theorem 1.1 to the unitary estimation algorithm due to Yang, Renner, and Chiribella [YRC20] (see also [Kah07, YYM25]).
@@
Corollary 1.4 is obtained by applying Theorem 1.1 to the state estimation algorithm due to Chen [Che25].
@@ 1.2 Discussion
While preparing this manuscript, we became aware of a concurrent work by Mele and Bittel [MB25], who independently obtained Corollary 1.2 via a different approach, including an explicit channel tomography algorithm. We also note that our Corollary 1.3 disproves a conjecture posed in [MB25], which conjectured that the upper bound $O(rd_{1}d_{2}/\varepsilon^{2})$ is optimal across the full parameter regime. This raises a new question: how does the true query complexity of quantum channel tomography behave near the boundary $rd_{2}\approx d_{1}$? We conjecture that the transition between classical and Heisenberg scalings is “smooth”, in the sense that the complexity exhibits a mixture of these two scalings.
@@ 2.2 Quantum channels
where $E_{i}:\mathbb{C}^{d_{1}}\rightarrow\mathbb{C}^{d_{2}}$ are non-zero linear operators that satisfy $\sum_{i=1}^{r}E_{i}^{\dagger}E_{i}=I$, which are called Kraus operators. We can always find a set of $E_{i}$ such that $\operatorname{tr}(E_{i}^{\dagger}E_{j})=0$ for $i\neq j$, then those $E_{i}$ are called orthogonal Kraus operators and $r$ is called the Kraus rank. In particular, a quantum channel that has Kraus rank $r=1$ is an isometry channels $\mathcal{V}=V(\cdot)V^{\dagger}$, where $V:\mathbb{C}^{d_{1}}\rightarrow\mathbb{C}^{d_{2}}$ is an isometry operator, i.e., $V^{\dagger}V=I_{d_{1}}$, and it must hold that $d_{1}\leq d_{2}$.
@@ 2.3 Formalism of quantum channel testers
Definition 2.4 (Link product $\star$[CDP09]).
@@ 2.4 Schur-Weyl duality on bipartite systems
$\bigotimes_{i=1}^{n}\mathcal{H}_{i}\stackrel{{\scriptstyle\mathfrak{S}_{n}\otimes\mathbb{U}_{d}}}{{\cong}}\bigoplus_{\lambda\vdash_{d}\,n}\mathcal{P}_{\lambda}\otimes\mathcal{Q}^{d}_{\lambda},$ (4)
@@ 3 Local test of quantum channel
Notation 3.2.
@@ Appendix A Isometry channel tomography in diamond norm
In this appendix, we prove the following lemma, which extends an algorithm for unitary channel tomography [HKOT23] to isometry channel tomography.
@@
Let $d_{1}\leq d_{2}$ be two positive integers, $V:\mathbb{C}^{d_{1}}\rightarrow\mathbb{C}^{d_{2}}$ be an isometry and $\mathcal{V}=V(\cdot)V^{\dagger}\in\textbf{{ISO}}_{d_{1},d_{2}}$ be the corresponding isometry channel. There exists an algorithm that uses $O(d_{1}d_{2}/\varepsilon^{2})$ queries to $\mathcal{V}$ and outputs an isometry channel estimate $\widehat{\mathcal{V}}$ such that $\|\mathcal{V}-\widehat{\mathcal{V}}\|_{\diamond}\leq\varepsilon<1$ with probability $\geq 2/3$. Moreover, these queries are used in parallel.
@@
with probability $\geq 1-2\eta$. By [AKN98, Lemma 12] and the triangle inequality,
$\left\lVert\mathcal{V}-\widehat{\mathcal{V}_{2}}\Phi^{\dagger}\mathcal{F}^{\dagger}\right\rVert_{\diamond}$ $\leq 2\left\lVert V\langle 0\rvert\Phi_{2}\lvert 0\rangle-\widehat{V_{2}}\Phi^{\dagger}F^{\dagger}\right\rVert_{\textup{op}}$
@@
$\leq 2\left\lVert V-\widehat{V_{2}}\Phi_{2}^{\dagger}F^{\dagger}\right\rVert_{\textup{op}}+2\left\lVert\langle 0\rvert\Phi_{2}\lvert 0\rangle\cdot\Phi-\Phi_{2}\right\rVert_{\textup{op}}$
$\leq 98\varepsilon,$
with probability $\geq 1-2\eta$, where the last inequality uses Equations 23 and 29. Finally, our algorithm outputs the isometry channel corresponding to $\widehat{V^{\prime}}=\widehat{V_{2}}\Phi^{\dagger}F^{\dagger}$ as the estimate. ∎
@@
by combining with Equations 34 and 35, we have

### v2

@@ 1.1 Our results
Theorem 1.1 can be viewed as an extension of the previous result by [CWZ24], which studies the power of local test for quantum states and shows that access to purifications does not help for mixed-state testing. Related results trace back to [SW22, Theorem 35], and were recently strengthened in an algorithmic sense by [TWZ25], which explicitly constructs an algorithm for generating random purifications of a mixed state (see also [GML25]). This is further leveraged in [PSTW25] for optimal mixed-state tomography. Intuitively, local test and random purification can be viewed as dual concepts in the Heisenberg and Schrödinger pictures, respectively. Finally, we note that Theorem 1.1 partially answers a conjecture from [TWZ25] asserting that access to channel dilations does not help.
@@
Notably, when $rd_{2}=d_{1}$ (i.e., when the quantum channel $\mathcal{E}$ can be obtained from a unitary channel followed by tracing out a subsystem), we can achieve the Heisenberg scaling $O(1/\varepsilon)$, even if $\mathcal{E}$ is not a unitary channel. We call the parameter regime $rd_{2}=d_{1}$ as the boundary regime, since any quantum channel must satisfy the constraint $rd_{2}\geq d_{1}$.
@@
Corollary 1.3 is obtained by applying Theorem 1.1 to the unitary estimation algorithm due to Yang, Renner, and Chiribella [YRC20], which achieves the Heisenberg scaling using parallel queries (see also [Kah07, YYM25]).
@@
Corollary 1.4 is obtained by applying Theorem 1.1 to the state estimation algorithm due to Chen [Che25], which achieves the Heisenberg scaling using parallel queries.
@@ 1.2 Discussion
While preparing this manuscript, we became aware of an independent and concurrent work by Mele and Bittel [MB25], who established the same upper bound $O(rd_{1}d_{2}/\varepsilon^{2})$ for quantum channel tomography in diamond norm error. They also provided an explicit and non-trivial dependence on the failure probability. We note that their method and ours are based on different techniques. They obtain the upper bound by analyzing tomography of Choi states, which yields an explicit tomography algorithm. In contrast, our approach is based on simulating access to dilations of quantum channels using the local test techniques, thereby reducing general channel tomography to a more tractable task—isometry channel tomography. In addition, we also provide upper bounds with Heisenberg scaling $O(1/\varepsilon)$ for (possibly non-unitary) channels in the boundary regime $rd_{2}=d_{1}$ (i.e., the regime in which channels admit unitary Stinespring dilations).
1.2.1 Open questions
Corollary 1.3 shows that the error dependence can be improved from the classical scaling $O(1/\varepsilon^{2})$ to the Heisenberg scaling $O(1/\varepsilon)$ in the boundary regime $rd_{2}=d_{1}$. This raises a new question: how does the true query complexity of quantum channel tomography behave in the near-boundary regime $rd_{2}\approx d_{1}$? We conjecture that the transition between classical and Heisenberg scalings is “smooth”, in the sense that the complexity exhibits a mixture of these two scalings.
Finally, we note that this paper shows dilations do not help for parallel testers, while the conjecture in [TWZ25] remains open in full generality: can one prove that dilations do not help for arbitrary (e.g., sequential) testers?
@@ 2.2 Quantum channels
where $E_{i}:\mathbb{C}^{d_{1}}\rightarrow\mathbb{C}^{d_{2}}$ are non-zero linear operators that satisfy $\sum_{i=1}^{r}E_{i}^{\dagger}E_{i}=I$, which are called Kraus operators. We can always find a set of $E_{i}$ such that $\operatorname{tr}(E_{i}^{\dagger}E_{j})=0$ for $i\neq j$, then those $E_{i}$ are called orthogonal Kraus operators and $r$ is called the Kraus rank. Note that $r$ must satisfy the constraint $d_{1}/d_{2}\leq r\leq d_{1}d_{2}$. In particular, a quantum channel that has Kraus rank $r=1$ is an isometry channels $\mathcal{V}=V(\cdot)V^{\dagger}$, where $V:\mathbb{C}^{d_{1}}\rightarrow\mathbb{C}^{d_{2}}$ is an isometry operator, i.e., $V^{\dagger}V=I_{d_{1}}$, and it must hold that $d_{1}\leq d_{2}$.
@@ 2.3 Formalism of quantum channel testers
Definition 2.4 (Link product $\star$[CDP08, CDP09]).
@@ 2.4 Schur-Weyl duality on bipartite systems
$\bigotimes_{i=1}^{n}\mathcal{H}_{i}\stackrel{{\scriptstyle\mathfrak{S}_{n}\times\mathbb{U}_{d}}}{{\cong}}\bigoplus_{\lambda\vdash_{d}\,n}\mathcal{P}_{\lambda}\otimes\mathcal{Q}^{d}_{\lambda},$ (4)
@@ 3 Local test of quantum channel
Notation 3.2 (Estimation tasks of quantum channels).
@@ Appendix A Isometry channel tomography in diamond norm
In this appendix, we prove the following lemma, which extends the $O(d^{2}/\varepsilon^{2})$ unitary channel tomography algorithm in [HKOT23] to isometry channel tomography.
@@
Let $d_{1}\leq d_{2}$ be two positive integers, $\varepsilon\in(0,1)$, $V:\mathbb{C}^{d_{1}}\rightarrow\mathbb{C}^{d_{2}}$ be an isometry and $\mathcal{V}=V(\cdot)V^{\dagger}\in\textbf{{ISO}}_{d_{1},d_{2}}$ be the corresponding isometry channel. There exists an algorithm that uses $O(d_{1}d_{2}/\varepsilon^{2})$ queries to $\mathcal{V}$ and outputs an isometry channel estimate $\widehat{\mathcal{V}}$ such that $\|\mathcal{V}-\widehat{\mathcal{V}}\|_{\diamond}\leq\varepsilon$ with probability $\geq 2/3$. Moreover, these queries are used in parallel.
@@
with probability $\geq 1-2\eta$. Then, we have
$\left\lVert\mathcal{V}-\widehat{\mathcal{V}_{2}}\Phi^{\dagger}\mathcal{F}^{\dagger}\right\rVert_{\diamond}$ $\leq 2\left\lVert V\langle 0\rvert\Phi_{2}\lvert 0\rangle-\widehat{V_{2}}\Phi^{\dagger}F^{\dagger}\right\rVert_{\textup{op}}$ (30)
@@
$=2\left\lVert V-\widehat{V_{2}}\Phi_{2}^{\dagger}F^{\dagger}\right\rVert_{\textup{op}}+2\left\lVert\langle 0\rvert\Phi_{2}\lvert 0\rangle\cdot\Phi-\Phi_{2}\right\rVert_{\textup{op}}$ (31)
$\leq 98\varepsilon,$ (32)
with probability $\geq 1-2\eta$, where Equation 30 is by [AKN98, Lemma 12] (see also [KSW08]), Equation 31 uses the fact that $\Phi_{2}$ is a diagonal unitary, $\widehat{V_{2}}$ is an isometry and $F$ is a unitary, and Equation 32 uses Equations 23 and 29. Finally, our algorithm outputs the isometry channel corresponding to $\widehat{V^{\prime}}=\widehat{V_{2}}\Phi^{\dagger}F^{\dagger}$ as the estimate. ∎
@@
by combining with Equations 37 and 38, we have